Find the vertices, eccentricity, foci and asymptotes of the hyperbola (x2/8) − (y2 /4) = 1. Also trace it. Under what conditions on λ the line x+λy = 2 will be tangent to this hyperbola? Explain geometrically.


1
Expert's answer
2021-06-08T09:36:08-0400
"\dfrac{x^2}{8}-\dfrac{y^2}{4}=1"

"c^2=a^2+b^2=8+4=12"


Vertices: "(-a, 0), (a, 0)"


Vertices: "(-2\sqrt{2}, 0), (2\sqrt{2}, 0)"


Eccentricity "\dfrac{c}{a}=\dfrac{2\sqrt{3}}{2\sqrt{2}}=\sqrt{\dfrac{3}{2}}=\dfrac{\sqrt{6}}{2}>1"


Foci: "(-c, 0), (c, 0)"


Foci: "(-2\sqrt{3}, 0), (2\sqrt{3}, 0)"


"\dfrac{(-x)^2}{8}-\dfrac{y^2}{4}=\dfrac{x^2}{8}-\dfrac{y^2}{4}"

The curve is symmetrical about the "y" -axis.



"\dfrac{x^2}{8}-\dfrac{(-y)^2}{4}=\dfrac{x^2}{8}-\dfrac{y^2}{4}"

The curve is symmetrical about the "x" -axis.


The curve is symmetrical in opposite quadrants.


The curve is not symmetrical about the line "y=x."


The origin does not lie on the curve.


"y" - axis: "x=0=>-\dfrac{y^2}{4}=1." There is no solution.

There is no "y" - intercept.


"x" - axis: "y=0=>\dfrac{x^2}{8}=1, x=\pm2\sqrt{2}."

"y" - intercepts: "(-2\sqrt{2}, 0), (2\sqrt{2}, 0)."


Center ": (0, 0)"

First asymptote: "y=-\dfrac{b}{a}x=-\dfrac{\sqrt{2}}{2}x"


Second asymptote: "y=\dfrac{b}{a}x=\dfrac{\sqrt{2}}{2}x"



First directrix: "x=-\dfrac{4\sqrt{3}}{3}"


Second directrix: "x=\dfrac{4\sqrt{3}}{3}"




"\dfrac{x^2}{8}-\dfrac{y^2}{4}=1"

If "\lambda =0," then the vertical line "x=2" is not tangent to this hyperbola, since the vertices are "(-2\sqrt{2}, 0), (2\sqrt{2}, 0)."


The equation of a line is represented by


"y=-\dfrac{1}{\lambda}x+\dfrac{2}{\lambda}, \lambda\not=0"

The slope of the tangent at the point "(x_1, y_1)" can be obtained by differentiating the equation of the hyperbola.

"\dfrac{2x}{8}-\dfrac{2yy'}{4}=0"

"x_1-2y_1y'|_{(x_1, y_1)}=0"

"y'|_{(x_1, y_1)}=\dfrac{x_1}{2y_1}"

The equation of the tangent line is


"y-y_1=\dfrac{x_1}{2y_1}(x-x_1)"

"y=\dfrac{x_1}{2y_1}x-\dfrac{x_1^2}{2y_1}+y_1"

"y=\dfrac{x_1}{2y_1}x-\dfrac{x_1^2-2y_1^2}{2y_1}"

"\dfrac{x_1^2}{8}-\dfrac{y_1^2}{4}=1"

"x_1^2-2y_1^2=8"

The equation of the tangent line is


"y=\dfrac{x_1}{2y_1}x-\dfrac{8}{2y_1}"

If "\dfrac{2}{\lambda}=-\dfrac{8}{2y_1}," then "\lambda=-\dfrac{y_1}{4}."


"\dfrac{x_1}{2y_1}=-\dfrac{4}{y_1}"

"x_1=-8"

"(-8)^2-2y_1^2=8"

"y_1=\pm2\sqrt{7}"


"Point (-8, -2\sqrt{7}), \lambda=\dfrac{\sqrt{7}}{4}"



"Point (-8, 2\sqrt{7}), \lambda=-\dfrac{\sqrt{7}}{4}"